OpenStudy (anonymous):

A penny is dropped from the top of the Chrysler Building (320m high). How fast is it moving when it hits the ground?

OpenStudy (mrnood):

Once again I suppose we have to ignore air resiustance. It seems to me that a question set on such a grand scale should explicitly state this. The formula for motion under constant acceleration (e.g. gravity) that you need in this case is v^2 = u^2 + 2 as v= final velocity u = initial velocity (= 0 in this case) a = acceleration of gravity (= 9.81 m/s^2) s = distance travelled (= 320m in this case) so put these into the equation and solve for v

OpenStudy (anonymous):

i got 25.01? but that is not the answer @MrNood

OpenStudy (mrnood):

No - that's not correct the value you want is v \[v ^{2}=u ^{2}+ 2 a s\] u = 0 a = 9.81 m/s^2 s= 320 m